Find consumer surplus. Given that pd=(6-Q) ², ps=14+Q

1
Expert's answer
2022-01-23T15:38:39-0500

"P_d=P_s,"

"36-12Q+Q^2=14+Q,"

"Q^2-13Q+22=0,"

"Q=2,~P=16,"

"Q=11,~P=25."